OpenStudy (anonymous):

Please Help!! The volume of a box(V) varies directly with its length(l). If a box in the group has a length of 30 inches, and the girth of 20 inches (perimeter of the side formed by the width and height), what is its height? Use k = 24. (Hint: Volume = length width height. Solve for length, and substitute into the equation for constant of proportionality.) ___ inches or ___ inches

OpenStudy (anonymous):

V=kl l = V/k 30 =V/24 V = 720 cu. in. height = V/[length*width] height = V/[30*width)